OpenStudy (anonymous):

Inverse of functions: g(x) =1/x + 2/3

terenzreignz (terenzreignz):

\[\Large g(x) = \frac1x + \frac23\] yes?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

bilis mag type puteek

terenzreignz (terenzreignz):

Well, usuals... You might want to skip to the essentials... replace g(x) with x and replace x with y. \[\Large x = \frac1y + \frac23\]

terenzreignz (terenzreignz):

Now... can you solve for y?

OpenStudy (anonymous):

find the lcd

terenzreignz (terenzreignz):

Sure... that might work, but it'll take a long time... why don't you transpose the 2/3 to the other side first? Remember, put all terms with y on one side, and all terms without, on the other.

OpenStudy (anonymous):

x-2/3=1/y

terenzreignz (terenzreignz):

That's good :) \[\Large x - \frac23 = \frac1y\] NOW you use lcd in the LEFT SIDE.

OpenStudy (anonymous):

baket 2/3 nilipat?? xDDD

terenzreignz (terenzreignz):

Kasi wala siyang y.

terenzreignz (terenzreignz):

I said... bring all terms without y to the other side.

OpenStudy (anonymous):

ai ou, xD ahahah na iimagine ko na y ung x. xD

terenzreignz (terenzreignz):

baka naman kailangan mo na magpahinga? XD

OpenStudy (anonymous):

hahahaha, xD sheet, may play pa kami sa eng taeee, cge continue na lang, xD

terenzreignz (terenzreignz):

oo nga continue... \[\Large x - \frac23= \color{red}?\] Find the LCD nga eh...

OpenStudy (anonymous):

hahah, xD 3x???

terenzreignz (terenzreignz):

... \[\large x = \frac{3x}{3}\] sige... tapos...?

OpenStudy (anonymous):

diba dapat 3-2x/3x =2/3????

terenzreignz (terenzreignz):

teka natatanga ako... \[\Large x + \frac23 = \frac{3x}3 +\frac23 = \color{red}?\]

terenzreignz (terenzreignz):

mali pa yung symbol... - yun hindi dapat +

OpenStudy (anonymous):

san nanggaling 3x/3???

OpenStudy (anonymous):

feeling ko kailangan ko nang matulog,,

OpenStudy (anonymous):

idk sorry

terenzreignz (terenzreignz):

para pareho yung denominator doon sa 2/3 tandaan...: \[\Large \frac{a}c+ \frac{b}c = \frac{a+b}c\] para isa na lang yung fraction bar

OpenStudy (anonymous):

geh, geh,, pre matutulog na lang, ako mukhang need koh nah, xD waaaa salamaat ulit,, xDDD